OpenStudy (anonymous):

Find the x-coordinate of the solution for this system of equations. 4x + 2y = -4 x = 2y + 9

jimthompson5910 (jim_thompson5910):

4x + 2y = -4 4(2y+9) + 2y = -4 8y+36 + 2y = -4 10y+36 = -4 10y = -4-36 10y = -40 y = -40/10 y = -4 ------------------------- x = 2y+9 x = 2(-4)+9 x = -8+9 x = 1 So the answer is x = 1